# Jarník-Besicovitch type theorems for semisimple algebraic groups

### v1

@@ 6. An upper bound for the Hausdorff dimension of $S_{\gamma}^{c}$
$a_{-t}E_{w_{i}}(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)a_{t}\cdot\mathcal{S}(L)$
@@
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))^{2}\cdot\mathcal{Q}_{t,L}\supset\bigcup_{S\in\mathcal{P}_{t,L}}S.$
Meanwhile, the number of boxes in $\mathcal{Q}_{t,L}$ is at most
$\frac{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(a_{-t}E_{w_{i}}(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)a_{t})\cdot\mathcal{S}(L))}{\exp(\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t}))}.$
@@
$\theta_{i}\leq L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}.$
Note that $\operatorname{diam}\mathcal{G}_{t}\sim e^{\nu_{0}(a_{-t})}$ and $\nu_{0}(a_{-t})\neq 0$ by Lemma 6.9. Define
@@
Then by definition and Proposition 6.10, the subset $\mathcal{G}^{k}$ is a cover of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$ for any $k\in\mathbb{N}$ and $\operatorname{diam}\mathcal{G}^{k}\sim e^{\nu_{0}(a_{-k})}$.
@@
$\lesssim$ $\sum_{t\in\mathbb{N}}\frac{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(a_{-t}E_{w_{i}}(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)a_{t})\cdot\mathcal{S}(L))}{\exp(\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t}))}\cdot e^{s\nu_{0}(a_{-t})}.$
@@
$\lesssim$ $\sum_{t\in\mathbb{N}}\sum_{\theta_{i}\leq L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}}(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)^{a_{w_{i}}+\epsilon}\cdot|\mathcal{F}(L)|$
$e^{\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\alpha(a_{-t})}e^{-\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t})}\cdot e^{s\nu_{0}(a_{-t})}.$
Fix $t\in\mathbb{N}$. Note that $\tau<\beta_{0}(w_{i}a_{-1}w_{i}^{-1})$. Choose $t_{0}\in\mathbb{Z}$ such that
$C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot t_{0}})^{\dim V_{\beta_{0}}}\leq\theta_{i}.$
Then by equation ($\ast$6) before Lemma 6.9, we have
$\sum_{\theta_{i}\leq L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}}|\mathcal{F}(L)|/L^{a_{w_{i}}+\epsilon}$
$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\sum_{C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}}\leq L\leq C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot(l+1)})^{\dim V_{\beta_{0}}}}|\mathcal{F}(L)|/L^{a_{w_{i}}+\epsilon}$
$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\frac{1}{(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{a_{w_{i}}+\epsilon}}\sum_{L\leq C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot(l+1)})^{\dim V_{\beta_{0}}}}|\mathcal{F}(L)|$
$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\frac{1}{(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{a_{w_{i}}+\epsilon}}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot(l+1)})^{\dim V_{\beta_{0}}})^{A_{w_{i}}+\epsilon}$
$\sim$ $\sum_{t_{0}\leq l\leq t-1}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{A_{w_{i}}-a_{w_{i}}}.$
Combining all these equations, we obtain that
$\sum_{B\in\mathcal{G}^{k}}\operatorname{diam}(B)^{s}$
$\lesssim$ $\sum_{t\in\mathbb{N}}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)t})^{\dim V_{\beta_{0}}})^{a_{w_{i}}+\epsilon}\cdot e^{\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\alpha(a_{-t})}e^{-\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t})}\cdot e^{s\nu_{0}(a_{-t})}$
$\cdot\sum_{t_{0}\leq l\leq t-1}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{A_{w_{i}}-a_{w_{i}}}.$
By computing the series above in the cases $A_{w_{i}}<a_{w_{i}}$, $A_{w_{i}}=a_{w_{i}}$ and $A_{w_{i}}>a_{w_{i}}$, one can conclude that the series above converges if
$s>\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot(\max\{A_{w_{i}},a_{w_{i}}\}+\epsilon)\cdot\dim V_{\beta_{0}}.$
This implies that for any $\epsilon>0$
$\dim_{H}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))\leq\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot(\max\{A_{w_{i}},a_{w_{i}}\}+\epsilon)\cdot\dim V_{\beta_{0}}.$
@@
$\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\leq\max_{i\in\bar{I}}\left\{\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w_{i}},a_{w_{i}}\}\cdot\dim V_{\beta_{0}}\right\}$
@@
$\dim_{H}S_{\tau}^{c}\leq\max_{i\in\bar{I}}\left\{\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w_{i}},a_{w_{i}}\}\cdot\dim V_{\beta_{0}}\right\}.$
This completes the proof of Theorem 1.4.
@@
In this section, we prove Theorems 1.61.7 and 1.8.

### v2

@@ 6. An upper bound for the Hausdorff dimension of $S_{\rho}(\psi)^{c}$
$a_{-t}E_{w}(C_{\delta}\cdot e^{\lambda_{w}(\epsilon)\cdot t}/L)a_{t}\cdot\mathcal{S}(L)$
@@
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))^{2}\cdot\bigcup_{S\in\mathcal{Q}_{t,L}}S\supset\bigcup_{S\in\mathcal{P}_{t,L}}S.$
Moreover, the number of boxes in $\mathcal{Q}_{t,L}$ is at most (up to a constant)
$\frac{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}((a_{-t}E_{w}(C_{\delta}\cdot e^{\lambda_{w}(\epsilon)\cdot t}/L)a_{t})\cdot\mathcal{S}(L))}{\delta^{\dim R_{u}(\overline{\mathbf{P}}_{0})}\cdot\exp(\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t}))}.$
@@
$\theta_{w}\leq L\leq e^{\lambda_{w}(\epsilon)\cdot t}.$
Note that $\operatorname{diam}\mathcal{G}_{t}\asymp e^{\nu_{0}(a_{-t})}$ and $\nu_{0}(a_{-1})\neq 0$. Define
@@
Then by definition and Proposition 6.9, the subset $\mathcal{G}^{k}$ is a cover of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))$ for any $k\in\mathbb{N}$ and $\operatorname{diam}\mathcal{G}^{k}\asymp e^{\nu_{0}(a_{-k})}$.
@@
$\lesssim$ $\sum_{t\in\mathbb{N}}\frac{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(a_{-t}E_{w}(C_{\delta}\cdot e^{\lambda_{w}(\epsilon)\cdot t}/L)a_{t})\cdot\mathcal{S}(L))}{\exp(\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t}))}\cdot e^{s\nu_{0}(a_{-t})}.$
@@
$\lesssim$ $\sum_{t\in\mathbb{N}}\sum_{\theta_{w}\leq L\leq e^{\lambda_{w}(\epsilon)\cdot t}}(e^{\lambda_{w}(\epsilon)\cdot t}/L)^{a_{w}+\epsilon}\cdot|\mathcal{F}(L)|\cdot e^{\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\alpha(a_{-t})}e^{-\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t})}\cdot e^{s\nu_{0}(a_{-t})}.$
Fix $t\in\mathbb{N}$. Note that $\tau\leq\beta_{0}(wa_{-1}w^{-1})$. Choose $t_{0}\in\mathbb{Z}$ such that $e^{\lambda_{w}(\epsilon)\cdot t_{0}}\leq\theta_{w}.$ Then by the definition of $A_{w}$, we have
$\sum_{\theta_{w}\leq L\leq e^{\lambda_{w}(\epsilon)\cdot t}}|\mathcal{F}(L)|/L^{a_{w}+\epsilon}$
$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\sum_{e^{\lambda_{w}(\epsilon)\cdot l}\leq L\leq e^{\lambda_{w}(\epsilon)\cdot(l+1)}}|\mathcal{F}(L)|/L^{a_{w}+\epsilon}$
$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\frac{1}{(e^{\lambda_{w}(\epsilon)\cdot l})^{a_{w}+\epsilon}}\sum_{L\leq e^{\lambda_{w}(\epsilon)\cdot(l+1)}}|\mathcal{F}(L)|$
$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\frac{1}{(e^{\lambda_{w}(\epsilon)\cdot l})^{a_{w}+\epsilon}}(e^{\lambda_{w}(\epsilon)\cdot(l+1)})^{A_{w}+\epsilon}$
$\asymp$ $\sum_{t_{0}\leq l\leq t-1}(e^{\lambda_{w}(\epsilon)\cdot l})^{A_{w}-a_{w}}.$
Combining all the inequalities above, we obtain
(4) $\sum_{B\in\mathcal{G}^{k}}\operatorname{diam}(B)^{s}$
$\lesssim$ $\sum_{t\in\mathbb{N}}(e^{\lambda_{w}(\epsilon)\cdot t})^{a_{w}+\epsilon}\cdot e^{\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\alpha(a_{-t})}e^{-\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t})}\cdot e^{s\nu_{0}(a_{-t})}\cdot\sum_{t_{0}\leq l\leq t-1}(e^{\lambda_{w}(\epsilon)\cdot l})^{A_{w}-a_{w}}.$
Computing the series (4) in the cases $A_{w}<a_{w}$, $A_{w}=a_{w}$ and $A_{w}>a_{w}$, one can conclude that the series (4) converges if
$s>\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{\lambda_{w}(\epsilon)}{\nu_{0}(a_{1})}\cdot(\max\{A_{w},a_{w}\}+\epsilon).$
This implies that for sufficiently small $\epsilon>0$
$\dim_{H}(S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))$
$\leq$ $\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{\lambda_{w}(\epsilon)}{\nu_{0}(a_{1})}\cdot(\max\{A_{w},a_{w}\}+\epsilon).$
@@
$\dim_{H}(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\leq\max_{\begin{subarray}{c}w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}\\ \beta_{0}(wa_{-1}w^{-1})\\ \geq\tau\end{subarray}}\left\{\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}\right\}.$
Proof of Theorem 1.5.
By the Bruhat decomposition of $\mathbf{G}$, we have
$\mathbf{G}(\mathbb{R})=\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{0}(\mathbb{R})\cdot w\cdot\mathbf{P}_{0}(\mathbb{R})=\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{0}(\mathbb{R})\cdot w\cdot R_{u}(\mathbf{P}_{0})(\mathbb{R}).$
For any $w\in{{}_{\mathbb{Q}}\mathcal{W}}$, define
$\mathbf{U}_{w}=(w^{-1}\cdot R_{u}(\overline{\mathbf{P}}_{0})\cdot w)\cap R_{u}(\mathbf{P}_{0})\textup{ and }\mathbf{Q}_{w}=(w^{-1}\cdot\mathbf{P}_{0}\cdot w)\cap R_{u}(\mathbf{P}_{0}).$
Then $R_{u}(\mathbf{P}_{0})=\mathbf{Q}_{w}\cdot\mathbf{U}_{w}$ and we have
$\mathbf{G}(\mathbb{R})=$ $\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{0}(\mathbb{R})\cdot w\cdot(\mathbf{Q}_{w}(\mathbb{R})\cdot\mathbf{U}_{w}(\mathbb{R}))$
$=$ $\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{0}(\mathbb{R})\cdot(w\cdot\mathbf{U}_{w}(\mathbb{R})\cdot w^{-1})\cdot w$
$=$ $\mathbf{P}_{0}(\mathbb{R})\cdot R_{u}(\overline{\mathbf{P}}_{0})\cdot{{}_{\mathbb{Q}}\mathcal{W}}.$
By definition, if an element $g\in\mathbf{G}(\mathbb{R})$ belongs to $S_{\rho}(\psi)$, then for any $h\in\mathbf{P}_{0}(\mathbb{R})$ and any $w\in\mathbf{G}(\mathbb{Q})$, $h\cdot g\cdot w$ also belongs to $S_{\rho}(\psi)$. Therefore, we have
$S_{\rho}(\psi)^{c}=\mathbf{P}_{0}(\mathbb{R})\cdot(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\cdot{{}_{\mathbb{Q}}\mathcal{W}}$
@@
$\dim_{H}S_{\rho}(\psi)^{c}=\dim\mathbf{P}_{0}+\dim_{H}(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\leq$ $\max_{\begin{subarray}{c}w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}\\ \beta_{0}(wa_{-1}w^{-1})\geq\tau\end{subarray}}\left\{\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}\right\}.$
This completes the proof of Theorem 1.5. ∎
@@
In this section, we derive Theorems 1.6 from Theorems 1.4 and 1.5, and then prove Theorems 1.7 and 1.8. Let $\mathbf{G}$ be a connected $\mathbb{Q}$-simple group. We assume that the highest weight $\beta_{0}$ in $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ is a multiple of $\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}{\alpha}$ and write
$\beta_{0}=\kappa\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}{\alpha}$
for some constant $\kappa<0$,
For any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, define
$\mathbf{F}^{w}=R_{u}(\mathbf{P}_{0})\cap w^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w,\quad\mathbf{H}^{w}=R_{u}(\mathbf{P}_{0})\cap w^{-1}\mathbf{P}_{\beta_{0}}w.$
Since $R_{u}(\overline{\mathbf{P}}_{0})$ and $R_{u}(\mathbf{P}_{0})$ contains all the $\mathbb{Q}$-roots of $\mathbf{G}$ (with multiplicities), by [8, Proposition 21.9], we have
$\mathbf{F}^{w}\cdot\mathbf{F}_{w}=w^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w,\quad(w\mathbf{F}^{w}w^{-1})\cdot(w\mathbf{F}_{w}w^{-1})=R_{u}(\overline{\mathbf{P}}_{\beta_{0}}).$
So by [55, Proposition 5.26], the Haar measure $\mu_{R_{u}(\overline{\mathbf{P}}_{\beta_{0}})}$ on $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ can be written as a product of the Haar measure $\mu_{w\mathbf{F}^{w}w^{-1}}$ on $w\mathbf{F}^{w}(\mathbb{R})w^{-1}$ and the Haar measure $\mu_{w\mathbf{F}_{w}w^{-1}}$ on $w\mathbf{F}_{w}(\mathbb{R})w^{-1}$.
Note that in the representation $\rho_{\beta_{0}}:\mathbf{G}\to\operatorname{\mathbf{GL}}(\bigwedge^{\dim V_{\beta_{0}}}V)$ (the $\dim V_{\beta_{0}}$-exterior product of $\rho$), the linear space spanned by
$e_{V_{\beta_{0}}}:=e_{1}\wedge e_{2}\wedge\cdots\wedge e_{\dim V_{\beta_{0}}}\in\bigwedge^{\dim V_{\beta_{0}}}V$
is the unique weight space of highest weight $\dim V_{\beta_{0}}\cdot\beta_{0}$ relative to $\mathbf{T}$. Let $W_{\beta_{0}}\subset\bigwedge^{\dim V_{\beta_{0}}}V$ be the irreducible sub-representation of $\mathbf{G}$ in $\rho_{\beta_{0}}$ such that $W_{\beta_{0}}$ contains $\mathbb{C}\cdot e_{V_{\beta_{0}}}$. Then $\Psi_{w}$ and $a_{w},b_{w},c_{w},A_{w}$ ($w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$), which are defined in §6, can be considered as morphisms and quantities defined in the sub-representation $W_{\beta_{0}}$.
Lemma 7.1.
For any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, we have $a_{w}\leq a_{e}$ and $A_{w}\leq A_{e}$.
Proof.
Let $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$. The inequality $A_{w}\leq A_{e}$ follows from the fact that $w\mathbf{F}_{w}w^{-1}\subseteq R_{u}(\overline{\mathbf{P}}_{\beta_{0}})=\mathbf{F}_{e}$.
Now let $B_{w\mathbf{F}^{w}w^{-1}}(1)$ be the open ball of radius one around the identity in $w\mathbf{F}^{w}(\mathbb{R})w^{-1}$. For any $R>0$, consider the subset
$B_{w\mathbf{F}^{w}w^{-1}}(1)\cdot(w\cdot\Psi_{w}^{-1}(B_{R})\cdot w^{-1})$
in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})=\mathbf{F}_{e}(\mathbb{R})$. We know that there exists a constant $C>0$ such that for any $f\in B_{w\mathbf{F}^{w}w^{-1}}(1)$
$\|\rho_{\beta_{0}}(f)\cdot v\|\leq C\|v\|\quad\left(\forall v\in\bigwedge^{\dim\beta_{0}}V\right).$
By the definitions of the morphisms $\Psi_{w}$ and $\Psi_{e}$, it implies that for any $R>0$
$B_{w\mathbf{F}^{w}w^{-1}}(1)\cdot(w\cdot\Psi_{w}^{-1}(B_{R})\cdot w^{-1})\subset\Psi_{e}^{-1}(B_{C\cdot R})$
and we have
$\mu_{w\mathbf{F}^{w}w^{-1}}(B_{w\mathbf{F}^{w}w^{-1}}(1))\cdot\mu_{\mathbf{F}_{w}}(\Psi_{w}^{-1}(B_{R}))$
$=$ $\mu_{w\mathbf{F}^{w}w^{-1}}(B_{w\mathbf{F}^{w}w^{-1}}(1))\cdot\mu_{w\mathbf{F}_{w}w^{-1}}(w\cdot\Psi_{w}^{-1}(B_{R})\cdot w^{-1})$
$\leq$ $\mu_{\mathbf{F}_{e}}(\Psi_{e}^{-1}(B_{C\cdot R})).$
Note that for any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$
$\mu_{\mathbf{F}_{w}}(\Psi_{w}^{-1}(B_{R}))\sim c_{w}R^{a_{w}}(\log R)^{b_{w}}\;(\textup{as }R\to\infty).$
So $a_{w}\leq a_{e}$. This completes the proof of the lemma. ∎
By the irreducible representation $\rho_{\beta_{0}}:\mathbf{G}\to\operatorname{\mathbf{GL}}(W_{\beta_{0}})$, the flag variety $\mathbf{G}/\mathbf{P}_{\beta_{0}}$ can be considered as a subvariety in $\mathbf{P}(W_{\beta_{0}})$ via
$g\mathbf{P}_{\beta_{0}}\mapsto\rho_{\beta_{0}}(g)(\mathbb{C}\cdot e_{V_{\beta_{0}}}).$
For any rational point $x\in(\mathbf{G}/\mathbf{P}_{\beta_{0}})(\mathbb{Q})\subset\mathbf{P}(W_{\beta_{0}})(\mathbb{Q})$, define the height of $x$ by
$H_{\beta_{0}}(x):=\|v\|$
where $v$ is the unique primitive integral vector (up to sign) in the rational line representing $x\in(\mathbf{G}/\mathbf{P}_{\beta_{0}})(\mathbb{Q})$ in $W_{\beta_{0}}$. Then we may count the number of rational points in $(\mathbf{G}/\mathbf{P}_{\beta_{0}})(\mathbb{Q})$ with the height function $\|\cdot\|_{\beta_{0}}$. By [63, Theorem 4], there exist $a_{0},b_{0},c_{0}>0$ such that
$\left|\{x\in(\mathbf{G}/\mathbf{P}_{\beta_{0}})(\mathbb{Q}):H_{\beta_{0}}(x)\leq R\}\right|\sim c_{0}\cdot R^{a_{0}}(\log R)^{b_{0}}.$
Moreover, we may compute that $a_{0}=1/(|\kappa|\cdot\dim V_{\beta_{0}})$.
Lemma 7.2.
For $e\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, we have
$a_{e}\leq 1/(|\kappa|\cdot\dim V_{\beta_{0}}),\quad A_{e}\leq 1/(|\kappa|\cdot\dim V_{\beta_{0}}).$
Proof.
For any $R>0$, let $B_{R}$ denote the ball of radius $R>0$ around the origin in $W_{\beta_{0}}$, and $\chi_{R}$ the characteristic function of $B_{R}$. Let $\Omega_{e}$ be a bounded fundamental domain of $\mathbf{F}_{e}(\mathbb{R})/\mathbf{F}_{e}(\mathbb{Z})$ in $\mathbf{F}_{e}(\mathbb{R})$. Then $\Omega_{e}$ is relatively compact. This implies that there exists $C>0$ such that for any $g\in\Omega_{e}$
$\|\rho_{\beta_{0}}(g)\cdot v\|\leq C\|v\|\quad(\forall v\in W_{\beta_{0}}).$
So we have
$\mu_{\mathbf{F}_{e}}(\Psi_{e}^{-1}(B_{R}))=$ $\int_{\mathbf{F}_{e}(\mathbb{R})}\chi_{R}(u\cdot e_{V_{\beta_{0}}})d\mu_{\mathbf{F}_{e}}(u)$
$=$ $\sum_{\gamma\in\mathbf{F}_{e}(\mathbb{Z})}\int_{\Omega_{e}}\chi_{R}(u\cdot\gamma\cdot e_{V_{\beta_{0}}})d\mu_{\mathbf{F}_{e}}(u)$
$\leq$ $\sum_{\gamma\in\mathbf{F}_{e}(\mathbb{Z})}\int_{\Omega_{e}}\chi_{C\cdot R}(\gamma\cdot e_{V_{\beta_{0}}})d\mu_{\mathbf{F}_{e}}(u)$
$=$ $\mu_{\mathbf{F}_{e}}(\Omega_{e})\cdot\sum_{\gamma\in\mathbf{F}_{e}(\mathbb{Z})}\chi_{C\cdot R}(\gamma\cdot e_{V_{\beta_{0}}})$
$=$ $\mu_{\mathbf{F}_{e}}(\Omega_{e})\cdot|\{\gamma\in\mathbf{F}_{e}(\mathbb{Z}):\|\gamma\cdot e_{V_{\beta_{0}}}\|\leq C\cdot R\}|$
$\leq$ $\mu_{\mathbf{F}_{e}}(\Omega_{e})\cdot\left|\{x\in(\mathbf{G}/\mathbf{P}_{\beta_{0}})(\mathbb{Q}):H_{\beta_{0}}(x)\leq C\cdot R\}\right|.$
Here we use the injectivity of the morphism $\Psi_{e}$. Note that for $e\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ we have
$\mu_{\mathbf{F}_{e}}(\Psi_{e}^{-1}(B_{R}))\sim c_{e}R^{a_{e}}(\log R)^{b_{e}}\;(\textup{as }R\to\infty).$
It implies that
$a_{e}\leq 1/(|\kappa|\cdot\dim V_{\beta_{0}}).$
Now by the discussion in §6, there exists a constant $C_{\beta_{0}}>0$ depending on $\Omega_{\beta_{0}}$ such that
$|\{g\in\Omega_{\beta_{0}}\cap w\mathbf{F}_{w}w^{-1}(\mathbb{Q}):0pt(g)\leq l\}|\leq\left|\{x\in(\mathbf{G}/\mathbf{P}_{\beta_{0}})(\mathbb{Q}):H_{\beta_{0}}(x)\leq C_{\beta_{0}}\cdot l\}\right|.$
By taking the limit as $l\to\infty$, we obtain
$A_{e}\leq 1/(|\kappa|\cdot\dim V_{\beta_{0}}).$
This completes the proof of the lemma. ∎